# From LLMs to MLLMs to Agents: A Survey of Emerging Paradigms in Jailbreak Attacks and Defenses within LLM Ecosystem

### v2

@@ VI-B4 Hybrid Strategy Defense
Currently, defense frameworks that integrate multiple strategies remain the most effective and robust solutions. Ding et al. [26] propose a novel defense strategy based on analyzing the priority of LLM prompt processing. Some studies reveal significant vulnerabilities in existing LLM defense mechanisms when handling rewritten and nested jailbreak prompts. To address this, they introduce the following defense scheme: 1. Implement a safety-first prompting mechanism that guides the model to prioritize response safety through preemptive security review; 2. Integrate a harmfulness classifier to accurately detect potential jailbreak prompts; 3. Employ supervised fine-tuning (SFT) techniques to enhance the model’s robustness against specific jailbreak scenarios. Technically, Ding et al. use the Perplexity Filter [113] as the core detection tool, setting the window size to 10 and adopting the maximum perplexity of prompt window slices in the harmful behavior dataset as the threshold. Perplexity calculation is based on the GPT-2 model. Additionally, they incorporate the RA-LLM [135] framework, experimentally determining the optimal parameters: dropout rate set to 0.3, candidate count to 5, and decision threshold to 0.2. Chan et al. [120] propose three detection and defense mechanisms against system message attacks, including inserting reference keys, using a second LLM as an evaluator, and introducing self-reminders. Reference keys serve to identify potential tampering, the evaluator LLM compares the original and current instructions to detect anomalies, and self-reminders prompt the assistant to follow the initially set instructions before user input. These methods effectively identify and resist system message attacks, ensuring the accuracy and reliability of virtual assistant responses.
Subsequently, Chen et al. [126] propose StruQ, a defense method designed to counter prompt injection attacks. Its core idea is to use structured queries to strictly distinguish prompts from data, preventing the LLM from mistakenly treating malicious instructions embedded in user data as valid inputs. Specifically, StruQ consists of two key components: 1. a secure frontend that processes prompts and user data structurally, using special delimiters to clearly separate the two and filter out potentially malicious characters exploitable by attackers; 2. structured instruction tuning, a specialized training method that teaches the LLM to follow only the prompt instructions while ignoring potential attack content in user data. Panda et al. [130] present two training-free defense methods: Self-Improvement and External-Improvement. Self-Improvement leverages the LLM’s reasoning ability to enhance safety through self-assessment and self-correction. Specifically, after generating an initial response, the model self-reviews the output for potential violations; if issues are detected, it adjusts the answer according to safety standards and iteratively optimizes itself to reduce attack success rates (ASR) while maintaining normal instruction execution and minimizing over-rejection. External-Improvement relies on an external aligned model or few-shot examples for defense. This method introduces a rigorously aligned LLM as an external reference to assist in detection and answer optimization, and employs few-shot prompt engineering to encourage the model to adhere more strictly to safety norms during responses.
@@
Recently, Varshney et al. [58] propose various LLM defense methods, including adding safety instructions, providing contextual examples, performing input and output self-checks, incorporating unsafe examples in instruction tuning, and introducing contextual knowledge. These approaches balance reducing unsafe responses and avoiding excessive defense on safe inputs, with contextual examples combined with safety instructions and moderate unsafe samples showing notable tuning effects. However, incorporating contextual knowledge may cause the model to generate harmful responses, leading to over-defense and thus requires cautious application. Zhang et al. [131] propose the PsySafe defense framework, which includes input defense, psychology-based defense, and role-based defense. Input defense mainly filters inputs using dangerous content detectors but with limited effectiveness. Psychology-based defense identifies and mitigates agents’ dark psychological states, effectively reducing risky behaviors. Role-based defense adjusts inter-agent role configurations to suppress collective dangerous behaviors among agents. These methods address security vulnerabilities in multi-agent systems comprehensively from both external input filtering and internal psychological state regulation perspectives.
In summary, current defense technologies are still in a stage of gradual development. In particular, preventing jailbreak attacks in the context of multimodal inputs and intelligent agent systems has become a critical challenge, especially when striving to maintain model efficiency, diversity, and flexibility. On the one hand, existing security mechanisms are often designed for unimodal or static inputs, making them less effective against cross-modal attack paths and complex interactive scenarios. On the other hand, due to their capabilities in task planning, tool invocation, and memory management, agent systems introduce new vulnerabilities stemming from their openness and extensibility. Therefore, future research should focus on advancing defense strategies toward earlier stages and building systematic protective frameworks. This may include embedding defense mechanisms during training, designing dynamic detection systems based on behavioral pattern recognition, and integrating multi-level, fine-grained security auditing modules to enhance the model’s overall robustness and defense capabilities in real-world applications.
@@ VII Related Work
In recent years, the rapid development of LLM technologies gives rise to numerous novel jailbreak attack and defense methods. Systematically summarizing and analyzing these methods not only helps to fully understand their underlying principles and evolutionary trends, but also lays a solid theoretical foundation for building more robust and efficient defense mechanisms.
To comprehensively grasp the essence and development trajectory of jailbreak attacks, different studies propose their own classification frameworks in response to the growing variety of attack methods. Yi et al. [23] categorize attack approaches based on the transparency of the target model into black-box and white-box attacks, and divide defense mechanisms into prompt-level and model-level strategies. Shayegani et al. [136] classify existing research into three types based on learning structures: text-only attacks, multimodal attacks, and attacks targeting complex systems such as federated learning or multi-agent systems. Ma et al. [54] summarize attack techniques as adversarial attacks, data poisoning, backdoor attacks, jailbreak and prompt injection attacks, energy-delay attacks, data and model extraction attacks, and emerging agent-specific threats. They also summarize the corresponding defense strategies for each type of attack. Esmradi et al. [137] explore two categories of attacks: those targeting the model itself and those targeting model applications. The former typically requires professional expertise and longer execution time, while the latter is more accessible to attackers. Geiping et al. [53] provide a comprehensive overview of the potential attack surfaces and targets of LLMs, and systematize various types of unintended behavior induction, such as deception, model control, denial of service, and data extraction. Rao et al. [52] classify jailbreak attack methods according to their intent into three categories: Information Leakage, Misaligned Content Generation, and Performance Degradation. They further discuss the challenges of jailbreak detection, especially in terms of effectiveness when facing known attack surfaces.
Although these studies propose relatively comprehensive strategies for attacks and defenses, limitations remain. These include insufficient attention to intelligent agents, a lack of detailed investigation into hybrid jailbreak methods and complex experimental setups, and difficulty in covering the latest developments in jailbreak research. To bridge these gaps, we review over 100 relevant studies and provide a more fine-grained classification of existing jailbreak attacks and defense techniques, further highlighting the relationships between them. In addition, we survey current evaluation metrics and jailbreak datasets to ensure a comprehensive understanding of the latest research progress in this field.
@@ VIII-A3 Limitations of Jailbreak and Defense Methods
(a) Generalization Limitations of Methods: 1. Jailbreak Generalization: Most existing jailbreak methods exhibit limited generalizability due to overly targeted technical approaches, resulting in weak transferability. Many studies focus on customized solutions for specific model architectures or attack scenarios. For example, adversarial prompt-based attacks often require fine-tuning based on the output patterns of the target model. Such highly customized methods tend to degrade significantly in effectiveness when model architectures are updated or application contexts change. 2. Defense Generalization: Current defense strategies remain highly fragmented when faced with diverse jailbreak attacks. Most mainstream defenses passively respond to specific attack patterns, such as detecting and filtering known adversarial examples or matching particular prompt templates. This “patch-based” defense approach lacks systematic theoretical support and often responds slowly to novel attack variants. More importantly, existing defense mechanisms are often deeply tied to specific LLMs, making it difficult to develop transferable and general defense techniques [146, 147].

### v3

@@ VI-B4 Hybrid Strategy Defense
Currently, defense frameworks that integrate multiple strategies remain the most effective and robust solutions. Ding et al. [26] propose a defense strategy based on analyzing the priority of LLM prompt processing. They address vulnerabilities in existing defenses against rewritten and nested jailbreak prompts by introducing: 1) a safety-first prompting mechanism to prioritize response safety, 2) a harmfulness classifier to detect jailbreak prompts, and 3) supervised fine-tuning (SFT) to enhance robustness. At the technical implementation level, they employ the Perplexity Filter [113] as the core detection tool, using the maximum perplexity of prompt window slices from the harmful behavior dataset as the decision threshold. Meanwhile, they introduce the RA-LLM framework [135] and determine the optimal parameter configuration through experiments. Chan et al. [120] propose three defense mechanisms against system message attacks: 1) inserting reference keys to identify tampering, 2) using a second LLM as an evaluator to detect anomalies, and 3) introducing self-reminders to ensure the assistant follows initial instructions. These methods effectively protect against system message attacks, ensuring response accuracy.
Subsequently, Chen et al. [126] propose StruQ, a defense against prompt injection attacks. StruQ uses structured queries to separate prompts from data, preventing the LLM from executing malicious instructions. It includes a secure frontend with delimiters and filters, and structured instruction tuning to ensure the LLM follows only valid prompt instructions while ignoring injected content. Panda et al. [130] present two training-free defenses: Self-Improvement and External-Improvement. Self-Improvement leverages the LLM’s reasoning for self-assessment and correction, aligning outputs with safety standards while minimizing over-rejection. External-Improvement uses an aligned external model or few-shot examples to detect and refine responses, ensuring compliance with safety norms.
@@
Recently, Varshney et al. [57] propose several LLM defense methods, including safety instructions, contextual examples, self-checks on inputs and outputs, unsafe examples in instruction tuning, and contextual knowledge integration. These methods balance reducing unsafe responses while avoiding over-defense on safe inputs. Contextual examples combined with safety instructions and moderate unsafe samples yield notable tuning effects. However, integrating contextual knowledge can sometimes lead to harmful responses, necessitating cautious application. Zhang et al. [131] introduce PsySafe, a defense framework combining input filtering, psychology-based mitigation, and role-based control. Input defense filters harmful content, psychology-based defense reduces agents’ risky behaviors by addressing dark psychological states, and role-based defense adjusts configurations to limit collective threats. Together, these methods provide a comprehensive security strategy for multi-agent systems, covering both external filtering and internal regulation.
In summary, current defense technologies are still at an early stage of development and face significant challenges in effectively mitigating jailbreak attacks, especially in scenarios involving multimodal inputs and intelligent agent systems. Many existing security mechanisms, originally designed for unimodal or static contexts, are inadequate for countering cross-modal attacks and handling the complexities of dynamic, multi-turn interactions. Furthermore, intelligent agents introduce additional security risks through their task planning, tool invocation, and memory retrieval capabilities, which expand the potential attack surface and complicate defense strategies. Future research should prioritize early-stage defenses and systematic protection frameworks, such as training-integrated safeguards, behavior-based dynamic detection, and multi-level security auditing to improve robustness in real-world applications.
@@ VII Related Work
In recent years, the rapid development of LLM technologies gives rise to numerous novel jailbreak attack and defense methods. Systematically summarizing and analyzing these methods not only helps to fully understand their underlying principles and evolutionary trends, but also lays a solid foundation for building more robust and efficient defense mechanisms.
To comprehensively grasp the essence and development trajectory of jailbreak attacks, different studies propose their own classification frameworks in response to the growing variety of attack methods. Yi et al. [23] categorize attack approaches based on the transparency of the target model into black-box and white-box attacks, and divide defense mechanisms into prompt-level and model-level strategies. Shayegani et al. [136] classify existing research into three types based on learning structures: text-only attacks, multimodal attacks, and attacks targeting complex systems like multi-agent systems. Ma et al. [54] summarize attack techniques as adversarial attacks, data poisoning, backdoor attacks, jailbreak and prompt injection attacks, energy-delay attacks, data and model extraction attacks, and emerging agent-specific threats. They also summarize the corresponding defense strategies for each type of attack. Esmradi et al. [137] explore two categories of attacks: those targeting the model itself and those targeting model applications. The former typically requires professional expertise and longer execution time, while the latter is more accessible to attackers. Geiping et al. [53] provide a comprehensive overview of the attack surfaces and attack objectives associated with LLMs, and systematize various ways to induce unintended behaviors, including deception, model manipulation, denial of service, and data extraction. Rao et al. [52] classify jailbreak attack methods according to their intent into three categories: Information Leakage, Misaligned Content Generation, and Performance Degradation. They further discuss the challenges of jailbreak detection, especially in terms of effectiveness when facing known attack surfaces.
Although these studies offer relatively comprehensive views for attacks and defenses, several limitations remain. These include insufficient attention to intelligent agents, a lack of detailed investigation into hybrid jailbreak methods and complex experimental setups, and difficulty in covering the latest developments in jailbreak research. To bridge these gaps, we review over 100 relevant studies and provide a more fine-grained classification of existing jailbreak attacks and defense techniques, further highlighting the relationships between them. In addition, we survey current evaluation metrics and jailbreak datasets to ensure a comprehensive understanding of the latest research progress in this field.
@@ VIII-A3 Limitations of Jailbreak and Defense Methods
(a) Generalization Limitations of Methods: 1. Jailbreak Generalization: Most existing jailbreak methods exhibit limited generalizability due to overly targeted technical approaches, resulting in weak transferability. Many studies focus on customized solutions for specific model architectures or attack scenarios. For example, adversarial prompt-based attacks often require fine-tuning based on the output patterns of the target model. Such highly customized methods tend to degrade significantly in effectiveness when model architectures are updated or application contexts change. 2. Defense Generalization: Existing defense strategies lack robustness and generalization, struggling to cope with the diversity of jailbreak attacks. Most mainstream defenses passively respond to specific attack patterns, such as detecting and filtering known adversarial examples or matching particular prompt templates. This “patch-based” defense approach lacks systematic theoretical support and often responds slowly to novel attack variants. More importantly, existing defense mechanisms are often deeply tied to specific LLMs, making it difficult to develop transferable and general defense techniques [146, 147].